Albert's father is twice his age. AND HE IS SIX TIMES THE AGE OF HIS LITTLE SON. THE IR AGES ADD UP TO 76 YEARS. HOW OLD IS EACH?
Answer by stanbon(75887)

Albert's father is twice his age. AND HE IS SIX TIMES THE AGE OF HIS LITTLE SON. THE IR AGES ADD UP TO 76 YEARS. HOW OLD IS EACH?
---------------
Equations:
f = 2S
S = 6s
---------
f + S + s = 76
-----
Substitute and solve for "f"
f + (f/2) + (1/6)(1/2)f = 76
-----
Multiply thru by 12 to get:
12f + 6f + f = 12*76
19f = 12*76
f = 48 years (father's age)
father's Son = f/2 = 24 years
father's grandson = S/6 = 4 years
--------------------------------------
Cheers,
Stan H.
===============